### v1

# Attitude Trajectories Stabilizable by Linear Control

Abstract
It is quite often claimed, and correctly so, that linear methods cannot achieve global stability results for attitude control, and conversely that nonlinear control is essential in order to achieve (almost) globally stable tracking of general attitude trajectories. On account of this definitive result, and also because of the existence of powerful nonlinear control techniques, there has been relatively very little work analyzing the limits and performance of linear attitude control. It is the purpose of this paper to provide a characterization of the stability achievable for one class of linear attitude control problems, namely those leading to a constant quaternion difference. In this paper, we analytically derive a critical error angle below which linearized dynamics lead to natural marginal stability for such a system, and above which the system is unstable. The dynamics are then used to derive a locally stable linear attitude controller whose performance is validated using simulations.
Keywords: Linear Attitude Control, Constant Quaternion Difference.
I Introduction
The attitude of a vehicle may be controlled to track a desired trajectory beginning from an (almost) arbitrary initial condition using a variety of (almost) globally stable nonlinear control algorithms, such as [1][2]. The controller in [1] uses a quaternion error variable to derive the controller, and this is extended in several works such as [3] to axisymmetric bodies, in [4] for neural network applications, and in [5] to account for modelled dissipation. An orthogonal matrix based representation of the attitude error is used in [6] to achieve similar stability results, and has been extended in several works such as [7] for robust control, [8] for output-feedback attitude control, [9] for flexibile bodies, and [2] for constrained systems.
Despite the known limitations of linear attitude control, its use is still widespread on account of the associated simplicity, and linear response to principal angle errors. For example, [10] used a linearization of the double integral form of attitude dynamics to regulate large initial errors. A controller that is linear in terms of the quaternion vector component is described in [11], which has the associated limitation of windup when the quaternion error angle is greater than 180 degrees. Linear PD and LQR methods for controlling a spacecraft’s Euler-angle attitude are compared in [12] leading to the conclusion that LQR methods lower the required control effort for similar attitude errors. In [13], the authors describe a linear attitude controller using Euler angle errors for a quadrotor UAV, and the controller works as long as the quadrotor’s attitude is sufficiently far away from the gimbal-lock singularity. In fact, such linear controllers are still prevalent in many commercial jet airplane autopilots which are for the most part far away from the singular Euler angle attitude. Other similar examples from literature include a design using the truncated Taylor series approximation for the quaternion exponentials in order to achieve a fast single-axis reorientation slew-rate on agile spacecrafts in [14], a PID attitude controller in [15], and an optimal linear controller in [16].
In this paper, we are interested in analyzing the stability of linear quaternion differential equations in time. In particular we consider the special case of a constant quaternion difference, since the results for this special case are sufficiently novel (the results have not been published previously in the literature to the best of the authors’ knowledge) and remarkable (being one of the few problems whose solutions can be analyzed and solved by hand). The chief contributions of this article may be summarized by the conditions provided on the error angle and the quaternion vector error in (14) and (15) for ensuring constant difference constant angle attitude tracking, the characteristic polynomial (28) for the linearized system, and the discriminant in (30) which determines marginal stability.
II Notation and Terminology
The attitude of a rigid body will be represented by a unit-magnitude 4-component quaternion $\check{q}=[q_{0}\quad q_{1}\quad q_{2}\quad q_{3}]^{T}$, with a check accent above the symbol representing it. The first component $q_{0}$ is also referred to as the scalar component, while the three components $[q_{1}\quad q_{2}\quad q_{3}]^{T}$ are also referred to as the vector part $\vec{q}_{v}$ of $\check{q}$. The most important law of quaternion algebra is their multiplication rule [17]:
$\check{p}\otimes\check{q}$ $=\begin{bmatrix}p_{0}\\ \vec{p}_{v}\end{bmatrix}\otimes\begin{bmatrix}q_{0}\\ \vec{q}_{v}\end{bmatrix}=\begin{bmatrix}p_{0}q_{0}-\vec{p}_{v}^{T}\vec{q}_{v}\\ p_{0}\vec{q}_{v}+q_{0}\vec{p}_{v}+\vec{p}_{v}\times\vec{q}_{v}\end{bmatrix},$ (1)
where $\vec{p}_{v}\times\vec{q}_{v}$ (third term in the vector part of the product) denotes the usual vector product of two Euclidean vectors in three-dimensional Euclidean space. The skew-symmetric matrix corresponding to the vector cross product will be denoted as $[\vec{p}_{v}\times]$ so that $\vec{p}_{v}\times\vec{q}_{v}$ may also be expressed in matrix notation as $[\vec{p}_{v}\times]\vec{q}_{v}$. It follows from the above multiplication rule that the multiplicative identity is the quaternion with unit scalar part and zero vector part, and that the inverse of a unit-magnitude quaternion is given by negating its vector part [17]:
$\check{q}$ $=\begin{bmatrix}q_{0}\\ \vec{q}_{v}\end{bmatrix}\Rightarrow\check{q}^{-1}=\begin{bmatrix}q_{0}\\ -\vec{q}_{v}\end{bmatrix}.$ (2)
We will refer to unit-magnitude quaternions simply as unit quaternions for brevity.
It may be recalled that unit quaternions provide a minimal, singularity-free, and global description of rotations [18], with composition of rotations achieved by multiplication of quaternions. The kinematics of an attitude quaternion may be expressed in terms of the body-frame angular velocity $\vec{\omega}$ as [17]
$\dot{\check{q}}$ $=\frac{1}{2}\check{q}\otimes\vec{\omega}.$ (3)
When Euclidean 3-vectors such as $\vec{\omega}$ are used as quaternion operands as in (3), they are prefixed with a scalar component of zero.
User control over rigid-body rotation is ultimately exercised not through the angular velocity itself, but rather through torques (or moments of forces) which influence the angular acceleration and hence the angular velocity. The relationship between the angular velocity and the torque may be described by Euler’s equations in vector form as
$\vec{n}$ $=\vec{\omega}\times J\vec{\omega}+J\dot{\vec{\omega}},$ (4)
where $\vec{n}\in\mathbb{R}^{3}$ is the net external torque, and $J\in\mathbb{R}^{3\times 3}$ is the (second) moment of inertia of the rigid body in the body-fixed coordinate frame. Assuming a solid body, the moment of inertia is positive definite, and the above equation may be inverted to obtain a one-to-one correspondence between the angular acceleration and the applied external torque. Therefore, one may justifiably consider the angular acceleration itself as the user input instead of the torque.
The problem we are concerned with in this article is that of controlling the attitude $\check{q}$ of a rigid body so as to track a reference attitude trajectory $\check{p}(t)$, whose kinematics are specified to be
$\dot{\check{p}}$ $=\frac{1}{2}\check{p}\otimes\vec{\nu},$ (5)
where $\vec{\nu}$ is the reference angular velocity. We will assume that the angular acceleration $\dot{\vec{\nu}}$ is also given to us. We will sometimes use the half-angular-velocities, $\vec{v}$ for $\vec{\nu}/2$ and $\vec{w}$ for $\vec{\omega}/2$, to avoid the factors of 2 and reduce notational clutter. In terms of $\vec{v}$ and $\vec{w}$, we have
$\dot{\check{p}}$ $=\check{p}\otimes\vec{v},\;\dot{\check{q}}=\check{q}\otimes\vec{w}.$ (6)
III Linear constant difference attitude control
In this section, we will consider the linear stability of the constant difference attitude control problem. The problem is interesting on its own from a theoretical standpoint. From a practical perspective, the problem manifests, for example, in the presence of an unknown disturbance that is quasistatic (constant, or very slowly varying relative to time scales relevant to the problem) referred to the body-frame. In the presence of such a disturbance, we cannot ensure convergence of the attitude difference to zero. Instead, it is its derivative alone which can be driven to zero.
From a purely geometric perspective, the constant difference constraint implies that the angle between the two attitudes is also a constant. Let
$\check{r}$ $=\check{q}-\check{p}$ (7)
denote the constant difference between the desired attitude $\check{p}$ and the actual attitude $\check{q}$. The difference quaternion $\check{r}$ is clearly not a unit quaternion in general, but that does not preclude linear control based upon it.
We are therefore interested in studying the linear (small perturbation) stability of solutions to the quaternion differential equation
$\dot{\check{q}}=\dot{\check{p}}.$ (8)
Define the error quaternion as
$\check{e}$ $=\check{p}^{-1}\otimes\check{q},$ (9)
with (6) yielding the dynamics
$\dot{\check{e}}$ $=-\vec{v}\otimes\check{p}^{-1}\otimes\check{q}+\check{p}^{-1}\otimes\check{q}\otimes\vec{w}$
$=\check{e}\otimes\vec{w}-\vec{v}\otimes\check{e}.$ (10)
The reader may recall our convention that pure vectors are prefixed with a scalar component of zero when used as quaternion operands. So $\vec{w}$ in $\check{e}\otimes\vec{w}$ is really the 4-vector $[0\quad\vec{w}^{T}]^{T}$.
The scalar part of the error rotation may be expressed in terms of $\check{r}$ and $\check{p}$ as
$e_{0}$ $=p_{0}q_{0}+\vec{p}_{v}^{T}\vec{q}_{v}=p_{0}(p_{0}+r_{0})+\vec{p}_{v}^{T}(\vec{p}_{v}+\vec{r}_{v})$
$=1+r_{0}p_{0}+\vec{r}_{v}^{T}\vec{p}_{v}.$ (11)
Similarly, $e_{0}$ may also be expressed in terms of $\check{r}$ and $\check{q}$ as
$e_{0}$ $=1-r_{0}q_{0}-\vec{r}_{v}^{T}\vec{q}_{v}.$ (12)
By adding up (11) and (12), and dividing by two, we obtain
$e_{0}$ $=1-|\check{r}|^{2}/2.$ (13)
Since the right hand side of the above equation is a constant, we see that $e_{0}$, and therefore the error angle, are also constant in time.
Using (6) in (8), we obtain the mathematical statement which encapsulates the constant difference constraint as a relationship between the desired and actual (half) angular velocities, $\vec{v}$ and $\vec{w}$:
$\vec{v}$ $=\check{e}\otimes\vec{w},$
or in terms of the scalar and vector parts,
$\vec{e}_{v}^{T}\vec{w}$ $=0,$ (14)
$\vec{v}$ $=e_{0}\vec{w}+\vec{e}_{v}\times\vec{w}.$ (15)
It may be noted that we could have as well used the constant difference constraint and the inverse of (9), to infer that $\vec{v}$ is also perpendicular to $\vec{e}_{v}$ like $\vec{w}$, and to express $\vec{w}$ in terms of $\vec{v}$ as follows
$\vec{w}=\check{e}^{-1}\otimes\vec{v}\;\Rightarrow\vec{e}_{v}^{T}\vec{v}=0,\;\vec{w}=e_{0}\vec{v}-\vec{e}_{v}\times\vec{v}.$ (16)
The scalar equation (14) implies that the body (half) angular velocity $\vec{w}$ must be perpendicular to the axis of the error rotation $\vec{e}_{v}$ in order to sustain a constant difference between $\check{p}$ and $\check{q}$. The relationship in (15) may be utilized to eliminate the desired (half) angular velocity $\vec{v}$ in (15) and express the kinematics of the error quaternion solely in terms of its own components $e_{0}$ and $\vec{e}_{v}$ and the actual (half) angular velocity $\vec{w}$:
$\dot{\check{e}}$ $=\begin{bmatrix}\vec{e}_{v}^{T}(\vec{v}-\vec{w})\\ e_{0}(\vec{w}-\vec{v})+\vec{e}_{v}\times(\vec{v}+\vec{w})\end{bmatrix}$
$=\begin{bmatrix}(e_{0}-1)\vec{e}_{v}^{T}\vec{w}\\ (e_{0}-1)\vec{w}+\vec{e}_{v}\times\vec{w}+\vec{e}_{v}\vec{e}_{v}^{T}\vec{w}\end{bmatrix},$ (17)
where we have expanded $\vec{v}$ in terms of $\vec{w}$ using (15).
We have so far utilized the relation in (15) to derive the kinematic governing equation (17). The scalar part of (17) simply yields the result derived earlier in (13), that the scalar component $e_{0}$ of the error quaternion, or equivalently the error angle, remains a constant when the angular velocity is perpendicular to the axis of the error rotation.
The vector part of (17) effectively contains three scalar equations in terms of six scalar components, three each in the vectors $\vec{e}_{v}$ and $\vec{w}$. In order to close the system, we need to bring in attitude kinetics and derive an equation for the body (half) angular acceleration $\dot{\vec{w}}$ which invokes the control torques and moments through (4).
While the kinematic equation (17) for $\dot{\vec{e}}_{v}$ cannot be simplified any further without additional assumptions, the kinetic equation allows us some liberty at attempting to achieve the constraint in (14). Taking the derivative of the scalar equation (14), and using (17), we obtain one scalar equation:
$\vec{e}_{v}^{T}\dot{\vec{w}}=-\vec{w}^{T}\dot{\vec{e}}_{v}=(1-e_{0})|\vec{w}|^{2}-(\vec{e}_{v}^{T}\vec{w})^{2}.$ (18)
It is worth emphasizing that although (18) does not explicitly invoke the desired angular acceleration $\dot{\vec{v}}$, the equation does in fact impose one scalar degree of constraint on $\dot{\vec{v}}$ on account of the relationship in (15). We may now design a minimal $\dot{\vec{w}}$ with no components along any other direction but $\vec{e}_{v}$:
$\dot{\vec{w}}$ $=\frac{(1-e_{0})|\vec{w}|^{2}\vec{e}_{v}}{|\vec{e}_{v}|^{2}}=\frac{|\vec{w}|^{2}\vec{e}_{v}}{1+e_{0}}.$ (19)
The two remaining degrees of freedom in $\dot{\vec{w}}$, namely the other two components along $\vec{w}$ and $\vec{e}_{v}\times\vec{w}$, may be passed on to designing $\dot{\vec{v}}$ without compromising what follows in this section. (See Section V for one possible linear design that yields asymptotic convergence on the angular velocity.)
Dynamical equations (17) and (19), together with the constant angle constraint $\dot{e}_{0}=0$, form the starting point for the following stability analysis. The governing equations for $\vec{e}_{v}$ and $\vec{w}$ may be written in state-space form as
$\begin{bmatrix}\dot{\vec{e}}_{v}\\ \dot{\vec{w}}\end{bmatrix}$ $=\begin{bmatrix}(e_{0}-1)\vec{w}+\vec{e}_{v}\vec{e}_{v}^{T}\vec{w}+\vec{e}_{v}\times\vec{w}\\ |\vec{w}|^{2}\vec{e}_{v}/(1+e_{0})\end{bmatrix}.$ (20)
We may linearize the above dynamics, and use the constant difference constraint $\vec{e}_{v}^{T}\vec{w}=0$, to obtain
$\begin{bmatrix}\dot{\tilde{e}}_{v}\\ \dot{\tilde{w}}\end{bmatrix}$ $=\begin{bmatrix}(e_{0}-1)\tilde{w}+\vec{e}_{v}\vec{w}^{T}\tilde{e}_{v}+\vec{e}_{v}\vec{e}_{v}^{T}\tilde{w}+\tilde{e}_{v}\times\vec{w}+\vec{e}_{v}\times\tilde{w}\\ (|\vec{w}|^{2}\tilde{e}_{v}+2\vec{e}_{v}\vec{w}^{T}\tilde{w})/(1+e_{0})\end{bmatrix}$
$=A\begin{bmatrix}\tilde{e}_{v}\\ \tilde{w}\end{bmatrix},$ (21)
where $\tilde{e}_{v}$ and $\tilde{w}$ are perturbations in the vector part $\vec{e}_{v}$ of the attitude error, and the (half) angular velocity $\vec{w}$, and
$A$ $=\begin{bmatrix}\vec{e}_{v}\vec{w}^{T}-[\vec{w}\times]&(e_{0}-1)1_{3\times 3}+[\vec{e}_{v}\times]+\vec{e}_{v}\vec{e}_{v}^{T}\\ \dfrac{|\vec{w}|^{2}1_{3\times 3}}{1+e_{0}}&\dfrac{2\vec{e}_{v}\vec{w}^{T}}{1+e_{0}}\end{bmatrix}.$ (22)
The trace of $A$ above may be immediately verified to be equal to zero after using (14). This implies that all the eigenvalues are on the imaginary axis, or at least one of them has a positive real part and the system is unstable.
Although the $6\times 6$ matrix $A$ in (22) looks quite formidable, it may be simplified significantly by applying the following orthogonal transformation. Let
$C$ $=\begin{bmatrix}\dfrac{\vec{e}_{v}}{|\vec{e}_{v}|}&\dfrac{\vec{w}\times\vec{e}_{v}}{|\vec{e}_{v}||\vec{w}|}&\dfrac{\vec{w}}{|\vec{w}|}\end{bmatrix};$ (23)
then,
$A^{\prime}=\begin{bmatrix}C^{T}\\ &C^{T}\end{bmatrix}A\begin{bmatrix}C\\ &C\end{bmatrix}$ (24)
[…]
where, $c=e_{0}$ and $s=|\vec{e}_{v}|$, hinting trigonometric ratios cosine and sine for the scalar and vector parts of the error quaternion $\check{e}$. For instance, $c$ would be the cosine of half the error angle. Note that empty entries in the matrix $A$ indicate a zero.
Since $A$ and $A^{\prime}$ are related through the orthogonal transformation in (24), they share the same characteristic polynomial. The characteristic polynomial $\det(\lambda 1_{6\times 6}-A^{\prime})$ for the matrix $A^{\prime}$ in (27) may be evaluated to be equal to
$f_{A}(\lambda)=f_{A^{\prime}}(\lambda)=\frac{\lambda^{6}}{|\vec{w}|^{6}}+\frac{(3-2e_{0}+e_{0}^{2})}{(1+e_{0})}\frac{\lambda^{4}}{|\vec{w}|^{4}}$
$\quad+\frac{(1-e_{0})(4+e_{0}+3e_{0}^{2})}{(1+e_{0})^{2}}\frac{\lambda^{2}}{|\vec{w}|^{2}}+\frac{2e_{0}^{2}(1-e_{0})^{2}}{(1+e_{0})^{3}}.$ (28)
An immediate observation is that the characteristic polynomial $f_{A}(\lambda)$ contains only even powers of $\lambda$ so roots are point symmetric about the origin. Multiplying $f_{A}$ with $(1+e_{0})^{3}$ and substituting $\lambda^{\prime}$ for $(1+e_{0})\lambda^{2}/|\vec{w}|^{2}$, $f_{A}(\lambda)$ may be simplified to the following cubic in terms of $\lambda^{\prime}$
$\lambda^{\prime 3}+(3-2e_{0}+e_{0}^{2})\lambda^{\prime 2}+(1-e_{0})(4+e_{0}+3e_{0}^{2})\lambda^{\prime}$
$\quad+2e_{0}^{2}(1-e_{0})^{2}.$ (29)
Since the roots of a cubic polynomial can be obtained in closed form, it is therefore possible to analytically compute the eigenvalues of the sixth order system in (21). This rather remarkable outcome is a consequence of the special structure associated with the dynamical equations (17) and (19) of the system under consideration (21), and their linearization about the constraint $\vec{e}_{v}^{T}\vec{w}=0$ in (14). In particular, the elements of $A$ may all be expressed in terms of standard Euclidean vector operations such as the scalar and vector product, and the projection operator. And the characteristics of systems described using such Euclidean vector expressions are invariant upon the use of orthogonal transformations. The specific orthogonal transformation $C$ used in (24) aligns the $x$-axis along $\vec{e}_{v}$ and the $z$-axis along $\vec{w}$.
The roots of the cubic in $\lambda^{\prime}$ are all real numbers when the non-trivial factor of the discriminant
$\!\!\!\!\Delta$ $=e_{0}^{6}+16e_{0}^{5}+70e_{0}^{4}+56e_{0}^{3}-151e_{0}^{2}+168e_{0}-112$ (30)
is positive (the trivial factors consist of twice repeated roots at $-1$ and $+1$). Within the range [1, 1]11[-1,\;1][ - 1 , 1 ] for $e_{0}$, the discriminant is positive when $e_{0}\gtrsim 0.85$ (error angle less than $\approx 31.7^{\circ}$) and negative otherwise. Furthermore, the coefficients in $f_{A}$ are nonnegative for all possible values of $e_{0}$, and the product of the coefficients of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient in (29), allowing us to conclude that the roots of (29) are always in the left half of the complex plane (that is, their real parts are negative). When they are purely real negative, that leads to purely imaginary roots in (28), and the system is marginally stable despite the lack of active negative feedback induced stabilization. When the roots of (29) are complex for $e_{0}\lesssim 0.84$ (error angle greater than $\approx 32.9^{\circ}$), that leads to (28) having roots with both positive and negative real parts and consequent instability.
The presence of purely imaginary eigenvalues for the system in (21) for bounded error angles tells us that the system is marginally stable and does not diverge away from its nominal difference, as long as the constraints in (14) and (15) are satisfied. This justifies designing a continuous switching controller comprising of a globally stable nonlinear controller for error angles above a certain bound, and a locally stable linear controller for small errors.
IV Validation using Simulations
In this section, we present simulation results to validate the key ideas presented in the previous section. The stability of the system is characterized by the eigenvalues of the state matrix in (22), which are the roots of the characteristic polynomial in (28). As we saw in the previous section, the best we can hope with respect to stability is that all the eigenvalues lie on the imaginary axis (since the trace of (22) is zero). In order that the $\lambda$ of (28) lie on the imaginary axis, the $\lambda^{\prime}$ of (29) must be negative real numbers. The latter condition may be verified by evaluating the discriminant (for realness) and applying the Routh-Hurwitz test (for negativity).
The first result is to demonstrate that all roots of (29) are in the left-half of the complex plane. This can be easily accomplished using the Routh-Hurwitz test (see Routh-Hurwitz stability criterion in, for example, [19]). Since $|e_{0}|\leq 1$, it is obvious that the coefficients of (29) are all non-negative. We also need to ensure that the product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient $c$. This condition is verified in Fig. 1.
Figure 1: The product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ in (29) must be greater than the constant coefficient $c$ in order to ensure all roots lie in the left half of the complex plane.
The next plot in Fig. 2 shows that the discriminant $\Delta$ of (29) changes its sign from positive to negative when $e_{0}\approx 0.84$. This marks the boundary of the unstable and marginally stable regions with respect to the error angle.
Figure 2: The discriminant of (29) may be plotted to graphically determine the critical error angle that separates the marginally stable region from the unstable region for the linearized dynamics in (21). Here, we see that the plot for the discriminant crosses the $x$-axis at $e_{0}$ approximately equal to 0.85, which prognosticates instability for values of $e_{0}$ less than approximately 0.84.
We now verify the cumulative outcome of the previous two plots, namely the condition that the eigenvalues $\lambda$ of (28) are in the left-half of the complex plane. Fig. 3 shows the locus of each of the six eigenvalues of the system matrix $A$ in (22). We can see that all six eigenvalues are indeed purely imaginary until $e_{0}$ falls below approximately 0.85. For $e_{0}$ less than approximately 0.84, the system has two purely imaginary eigenvalues, and four complex eigenvalues. Of the latter four, two lie in the positive half of the complex plane and hence indicate instability.
Figure 3: The root locus for the system matrix in (22) shows the movement of all six eigenvalues of the state matrix in. We can see that the linearized system whose dynamics are governed by (21) is marginally stable when $e_{0}\gtrsim 0.85$.
V Linear Attitude Control Design
We will now design a linear attitude controller using the marginally stable natural dynamics in (21) as a starting point, and by including feedback terms from the measured or estimated attitude and angular velocity errors, to move the eigenvalues of the system into the left half of the complex plane. To this end, we include four gain parameters, $K_{1}$ to $K_{4}$, in the feedback law as follows:
$\begin{bmatrix}\dot{\tilde{e}}_{v}\\ \dot{\tilde{\omega}}\end{bmatrix}$ $=A_{fb}\begin{bmatrix}\tilde{e}_{v}\\ \tilde{\omega}\end{bmatrix}$
where the feedback matrix $A_{fb}$ may be expressed in terms of the matrix $A$ in (22) as
$A+\begin{bmatrix}0&0\\ K_{1}1_{3\times 3}+K_{2}[w\times]&K_{3}1_{3\times 3}+K_{4}[e_{v}\times]\end{bmatrix}.$ (31)
The gains $K_{1}$ and $K_{3}$ provide the usual proportional and derivative feedback from the errors, while the gains $K_{2}$ and $K_{4}$ are introduced to take care of the cross-product terms in the error dynamics described in (20). Since we are considering linear control, $e_{v}$ and $w$ in (31) would be the nominal attitude error and the nominal body-frame (half) angular velocity.
Figure 4: The eigenvalues of the linearized attitude controller may all be moved into the left half of the complex plane by actively using negative feedback. The plots above correspond to a quaternion error of $0.11-e_{0}=$ 0.1, and use tuned feedback gains of $K_{1}=K_{3}=-0.3$, $K_{2}=1$, and $K_{4}=-0.5$.
The gains in (31) may now be tuned using simulations to stabilize the system by locating the eigenvalues in the left half of the complex plane resulting in the following values: $K_{1}=K_{3}=-0.3$, $K_{2}=1$, and $K_{4}=-0.5$. The root locus of the system, for feedback gains scaled by a factor varying from a tenth to ten times, is shown in Fig. 4.
VI Conclusion
We have looked at the limitations of linear attitude control at the kinematic level. One could also consider linear attitude control at the kinetic level, which would yield a second order differential equation. Some preliminary work suggests a constrained reference tajectory, similar to (14) and (15) for the kinematic equation, but a more detailed analysis might provide more insight into that problem.